OpenStudy (anonymous):

HELP!!!!!!! SOLVE: 4^x - 2= 8^6

OpenStudy (ash2326):

is this your question? \[4^{x-2}=8^6\]

OpenStudy (anonymous):

looks like 2 is not part of the power.....

OpenStudy (ash2326):

I assume it's that we have \(8^6\), it can be written as \[(8^2)^3= 64^3\] we know that \(4^3=64\) so \[(8^2)^3= 64^3=(4^3)^3\] this is equal to \(4^9\) so we have \[4^{x-2}=4^9\] equating the powers we get \[x-2=9\] add 2 to both sides, we get \[x=11\]

OpenStudy (anonymous):

Or, You can simply take it in the base of two. \[4^{x-2} = 8^{6}\] \[(2^{2})^{x-2} = (2^{3})^{6}\] Therefore we have, 2x - 4 = 18 => x = 11. Basically the aim is to get a same base. Irrespective of how you do it. :)

OpenStudy (anonymous):

AAHH YOU GUYS THANK YOU SOOO MUCH :3